# Possible principles for aligned structure learning agents

### v1

@@ Abstract
This paper offers a roadmap for the development of scalable aligned artificial intelligence (AI) from first principle descriptions of natural intelligence. In brief, a possible path toward scalable aligned AI rests upon enabling artificial agents to learn a good model of the world that includes a good model of our preferences. For this, the main objective is creating agents that learn to represent the world and other agents’ world models; a problem that falls under structure learning (a.k.a. causal representation learning). We expose the structure learning and alignment problems with this goal in mind, as well as principles to guide us forward, synthesizing various ideas across mathematics, statistics, and cognitive science. 1) We discuss the essential role of core knowledge, information geometry and model reduction in structure learning, and suggest core structural modules to learn a wide range of naturalistic worlds. 2) We outline a way toward aligned agents through structure learning and theory of mind. As an illustrative example, we mathematically sketch Asimov’s Laws of Robotics, which prescribe agents to act cautiously to minimize the ill-being of other agents. We supplement this example by proposing refined approaches to alignment. These observations may guide the development of artificial intelligence in helping to scale existing—or design new—aligned structure learning systems.
Keywords: Bayesian, generative model, world model, causal representation learning, active inference.
@@ 1 Introduction
We aim to be inclusive about—and relevant to—naturalistic approaches to artificial intelligence. For this we commit to a ‘first principles’ approach to modeling intelligence known as active inference [1, 2, 3]. Active inference is not divorced from other naturalistic approaches to modeling intelligence, but rather aims to accommodate them within a broader framework. Active inference follows a long lineage of ideas, perhaps originating from Helmholtz’s motion of perception as unconscious inference, which was reincarnated in the neurosciences at the turn of this century as predictive coding, and generalized as the Bayesian brain hypothesis. Active inference was proposed shortly afterward, in the mid 2000s, extending these Bayesian accounts by postulating that action optimizes the same objective as perception and learning [4, 5]. This account was suggested to be a potentially unifying brain theory, in the sense of being able to accommodate a wide range of previously existing and partially non-overlapping brain theories as special cases [6]. In light of the descriptive power of these ideas, researchers have sought to justify this account in terms of statistical physics, with increasing mathematical rigor and sophistication. These efforts have birthed a nascent field of non-equilibrium physics, known as ‘Bayesian mechanics’, that bridges stochastic descriptions of particles with inferential ones. This has been used to derive the active inference framework as we present it here, offering a description of sentient behaviour from first principles [7]. The active inference framework can be used to model a remarkable range of phenomena in cognitive science, ranging from human choice behavior [8] to psychopathology [9], to many known features of the brain’s anatomy and physiology [10, 11], including the activity of neural populations [12, 13]—see [1, 14, 2, 3] for reviews. Active inference has more recently gained traction in machine learning and robotics—see [15, 16, 17] for reviews of opportunities and challenges.
Learning generative (world) models:
@@
The key challenge to unlock the usefulness of this naturalistic approach to AI at scale is to enable agents to learn their model of the world, as current approaches to address this remain limited (but see [18, 19]). Note that this problem is not unique to active inference: it is shared across all of model-based reinforcement learning [20]. This structure learning problem [^1] Here we use the term structure learning synonymously with causal representation learning [21, 22] as is common in the neural and cognitive sciences [23]. That is, we use structure learning to denote the joint problem of learning causal relationships and latent representations. forms the focus of this article: we examine how AI systems and agents might tractably learn models of their data generating process. We examine this problem in detail and discuss the essentials roles of core knowledge, information geometry and model reduction, and suggest core structural modules to enable learning a wide range of naturalistic worlds.
@@ 2 First principles approach to natural intelligence
We aim for a ‘first principles’ description of natural intelligence. To this aim, we summarize a physical theory describing the dynamics of things that actively interact with their surrounding world.
@@
with potential function $\operatorname{G}$ known as expected free energy by analogy with the variational free energy over trajectories, and where $d\triangleq d_{t}\subseteq\{o_{\leq t},a_{\leq t},\mu_{\leq t}\}$ is a potentially non-empty history of data. Interestingly, (3) can be related to several information theoretic formulations of intelligent decision-making that predominate in statistics, cognitive science and engineering: see [31, 32, 33]. Please see[7, 26, 31, 34] for derivations of (3).
@@ 3.3 The prior: information geometry
Looking forward, we should step back and consider the natural problem of Bayesian inference on the space of models endowed with an information geometry, so that models that include the exact same information are identified, and so that we infer equivalence classes of models up to information invariance. This would prevent us from ‘over-counting’ models, which might be fundamental for structure learning at scale. Indeed, the number of symmetries that render two models equivalent increases vastly in the number of latent variables considered. These information invariances are not generally factored into current methods, which means that the model hypothesis space—hese methods face—is overwhelmingly larger than it should be. Quotienting by these (and other) invariances should greatly reduce the complexity of the problem (and highly improve model evidence). Furthermore, specifying the prior over equivalence classes directly has an advantage: considering that the prior probability of a model equivalence class equals the sum of the prior probabilities for all elements in the class, when specifying prior probabilities over individual models as opposed to equivalence classes, we may have unintended over-counting effects; leading to a higher prior for models with a high number of latent variables. Practically considering the problem in this way should be of great help for scalable structure learning.
@@ 4.3.1 Core knowledge priors
Core knowledge therefore constitutes knowledge that is valid across any naturalistic world, which translates into significant constraints on the prior over models. The prior over models as explanations for the world is then constrained by the consistency with the underlying information geometry (i.e. local constraints) and core knowledge constraints (i.e. non-local constraints). [^4] The maximal entropy probability distribution that is consistent with an underlying geometry and some constraints is a Gibbs measure on the underlying metric space [90, Chapter 9].
@@ 5 AI alignment
In what follows, we approach AI alignment through the lens of active inference and structure learning, using Asimov’s Three Laws of robotics [114] as a simple illustrative example. Our focus is on providing new ways of thinking about the AI alignment problem, rather than recommending a specific solution to be implemented. One conceptual point is that we can frame alignment as taking actions that comply with the other’s preferences, and we can learn those preferences through structure learning, which, in terms from psychology, corresponds to a sophisticated form of theory of mind.
5.1 Related work
The approach presented here shares important connections with other causal modeling approaches to AI alignment. Like us, the work of Everitt and colleagues uses causal models of agent-environment interactions, but with the aim of identifying agent’s incentives [115, 116]. This provides a complementary perspective on analyzing and designing AI systems with desirable incentive structures, and aligns well with our discussion of theory of mind, including its potential for benevolent as well as adversarial uses. The structure learning approach that we develop may provide a way to dynamically construct these causal networks, and go beyond identifying the mere presence or absence of various properties such as incentives [115], intentions [117], and deception [118], by quantifying these phenomena.
@@ 5.2 Well-being, alignment and cautious AI
where $d\triangleq d(t)$ is the data entertained by the agent at time $t$; necessarily a subset of past and present states of the agent (boundary and internal states). Note that well-being, when expressed in this way, is quantified in terms of natural units of information (nats). This definition of well-being is fairly established in the active inference literature [120, 121, 122, 123]. It is also possible to reproduce some well-known empirical phenomena of biological collectives subject to harm by quantifying harm in this way and following the active inference equations [124].
@@
Beyond the conceptual issues discussed here, active inference also offers algorithmic advances for building safer and more robust AI systems [16, 15].
@@ 5.3 Toward empathetic AI
We now define an empathetic agent as an agent that not only models the objects in its external world, but also the other agents within it—as well as their mental (i.e. internal) states. This effectively recognizes that the external world comprises other agents’ actions, observations and internal states, which may be inferred from data, along with all other external states in the shared environment. Technically, an agent that has beliefs about another’s mental state is capable of empathy. In the language of cognitive psychology, empathetic AIs have theory of mind.
Instantiating this in practice is a difficult problem, and may require entertaining generative world models as in Figure 10 (left panel). An empathetic agent may, in principle, infer another agent’s world model from the other’s inferred observations and actions (assuming, for instance, that it takes actions that minimize expected free energy—or any other objective—and solving the inverse problem). Note that assuming that the other possesses the same generative model as oneself ("you are like me") greatly simplifies the process by which one forms beliefs about another, because my model of me becomes my model of you [126, 127].
@@
However, it is crucial to recognize that empathy alone does not guarantee benevolence or safety. Competitive or even adversarial agents may benefit from sophisticated models of others for instrumental reasons. In particular they may use these to manipulate, or deceive others to achieve their goals [130, 131]. The key might be to prescribe agent’s actions to be benevolent in the sense of complying with the other’s preferences. However, note that benevolent agents may still choose to deceive, as it may happen to be beneficial to manipulate humans in the process of optimizing their well-being (e.g. white lies).
@@ 5.4 The Three Laws of Robotics
To mathematically instantiate the First Law, first consider an empathetic agent whose sole preferences are to prevent another agent (the target agent) from coming to harm. The empathetic agent’s actions could be governed by the expected free energy (4) where we discard all external variables besides the target’s harm
@@
A version of the First Law could be achieved by, e.g. assuming a preferred distribution over the target’s harm in (11) to be an exponentially decaying probability distribution over harm, so that zero or low harm would exponentially preferred to high harm, or by (soft-)thresholding the target’s well-being to desirable values. [^8] In any case, the dependence on the history of data in the preferred distribution means that the parameters of this distribution could be learned from data [36, 34]. To appropriately model the target’s level of harm, the empathetic agent would have to model many other external variables; however the only imperatives driving its actions would be to resolve ambiguity—and realize its preferences—about the target’s well-being.
@@ 5.5 Alignment beyond perspective taking
Numerical studies suggest that populations may naturally converge to free energy equilibria, where all agents share the same generative model (and preferences), albeit possibly at an evolutionary timescale [126, 133, 127]. In this case, each agent’s goal is the group’s (shared) goal and empathy is an emergent property (Section 5.3). Clearly, these numerical studies need to be extended to agents that do structure learning and learn their own preferences [36, 34].
@@ 6 Discussion
Active inference can be arrived at from two opposite directions: the bottom-up, inductive and historical approach, which is where theories of brain function were continuously refined and generalized to account for various empirical phenomena [134, 135, 5, 6]. A top-down, deductive approach, through the nascent field of physics called Bayesian mechanics [136, 27, 7], which bridges elementary descriptions of particles, things and agents in the physical world, with descriptions of inference. From a theoretical perspective, note that there are many opportunities to further develop Bayesian mechanics and the foundations of active inference.

### v2

@@ Abstract
This paper offers a roadmap for the development of scalable aligned artificial intelligence (AI) from first principle descriptions of natural intelligence. In brief, a possible path toward scalable aligned AI rests upon enabling artificial agents to learn a good model of the world that includes a good model of our preferences. For this, the main objective is creating agents that learn to represent the world and other agents’ world models; a problem that falls under structure learning (a.k.a. causal representation learning or model discovery). We expose the structure learning and alignment problems with this goal in mind, as well as principles to guide us forward, synthesizing various ideas across mathematics, statistics, and cognitive science. 1) We discuss the essential role of core knowledge, information geometry and model reduction in structure learning, and suggest core structural modules to learn a wide range of naturalistic worlds. 2) We outline a way toward aligned agents through structure learning and theory of mind. As an illustrative example, we mathematically sketch Asimov’s Laws of Robotics, which prescribe agents to act cautiously to minimize the ill-being of other agents. We supplement this example by proposing refined approaches to alignment. These observations may guide the development of artificial intelligence in helping to scale existing—or design new—aligned structure learning systems.
Keywords: Agent, world model, generative, model discovery, causal representation learning, Bayesian inference.
@@ 1 Introduction
We aim to be inclusive about—and relevant to—all naturalistic approaches to artificial intelligence. We approach this with a ‘first principles’ approach to modeling intelligence known as the active inference framework [1, 2, 3]. Active inference is not divorced from other naturalistic approaches to modeling intelligence, but rather aims to accommodate them within a broad framework derived from statistical physics. This follows a long lineage of ideas, perhaps originating from Helmholtz’s motion of perception as unconscious inference [4], which was reincarnated in the neurosciences at the turn of this century as predictive coding [5], and generalized as the Bayesian brain hypothesis [6]. Active inference was proposed shortly afterward, in the mid 2000s, extending these Bayesian accounts by postulating that action optimizes the same objective as perception and learning [7, 8]. This account was suggested to be a potentially unifying brain theory, in the sense of accommodating a wide range of previously existing and partially non-overlapping brain theories as special cases [9]. In light of the descriptive power of these ideas, researchers have sought to justify this account in terms of statistical physics, with increasing mathematical rigor and sophistication. These efforts have birthed a nascent field of non-equilibrium physics, known as ‘Bayesian mechanics’, that bridges stochastic descriptions of particles with inferential ones. This has been used to derive the active inference framework as we present it here, which provides a description of sentient behaviour [10]. The active inference framework can be used to model a remarkable range of phenomena in cognitive science, ranging from human choice behavior [11] to psychopathology [12], to many known features of the brain’s anatomy and physiology [13, 14], including the activity of neural populations [15, 16] (see [1, 17, 2, 3] for reviews). Active inference has more recently gained traction in machine learning and robotics (see [18, 19, 20] for reviews).
Learning world models:
@@
The key challenge to unlock the usefulness of model-based approaches to AI at scale is to enable agents to learn their model of the world, as current approaches to address this remain limited (but see [21, 22, 23]). Note that this problem is shared by active inference, model-based reinforcement learning and control [24]. This structure learning problem [^1] Here we use the term structure learning synonymously with causal representation learning [25, 26], or causal model discovery, as is common in the neural and cognitive sciences [27]. That is, we use structure learning to denote the joint problem of learning causal relationships and latent representations. forms the focus of this article: we examine how AI systems and agents might tractably learn models of their data generating process. We examine this problem in detail and discuss the essential roles of core knowledge, information geometry and model reduction, and suggest core structural modules to enable learning a wide range of naturalistic worlds.
@@
Our contribution:
This article provides a research roadmap toward scalable, aligned AI agents, identifying key research questions and promising directions of travel. This roadmap comprises three core elements: 1) A commitment to naturalistic approaches to AI that draw from biological intelligence. 2) An integrated framework for agents to autonomously learn world models, enabling scalable intelligence. This synthesizes insights from currently disparate fields—including machine learning, cognitive science, mathematics, and statistics. 3) A conceptual exploration of how these advances might address AI alignment challenges. In summary, this paper provides an integrative perspective and a point of reference for subsequent numerical studies.
Significance of framework:
While the roadmap we present is high-level and conceptual, recent work demonstrates the practical viability of key components. For instance, AXIOM [21] implements the structure learning principles outlined here (with object-centric active inference models) to master arcade games within 10,000 interaction steps—achieving remarkable sample, computational and parameter efficiency compared to deep reinforcement learning approaches. A similar line of work, theory-based reinforcement learning [28, 29] demonstrates that agents maintaining beliefs over probabilistic programs (encoding causal structure and parameters) can achieve human-level learning efficiency. These implementations illustrate core aspects of this roadmap and demonstrate that naturalistic, structure-learning approaches can deliver both interpretability and performance advantages over conventional methods. Our work further suggests that these and similar advances could be directly relevant to addressing AI alignment challenges.
@@ 2 First principles approach to natural intelligence
We aim for approaches to AI that build upon natural intelligence. In what follows, we outline a ‘first principles’ description of intelligent systems that is consistent with the natural sciences, particularly with physical descriptions of the natural world. Specifically, we present a physical theory describing the dynamics of systems that actively interact with their environment.
@@
with potential function $\operatorname{G}$ known as expected free energy by analogy with the variational free energy over trajectories, and where $d\triangleq d_{t}\subseteq\{o_{\leq t},a_{\leq t},\mu_{\leq t}\}$ is a potentially non-empty history of data. Interestingly, (3) can be related to several information theoretic formulations of intelligent decision-making that predominate in statistics, cognitive science and engineering: see [37, 38, 39]. See [10, 32, 37, 40] for derivations of (3).
@@ 3.3 The prior: information geometry
Looking forward, we should step back and consider the problem of Bayesian inference on the space of models endowed with an information geometry, so that models that include the exact same information are identified, and so that we infer equivalence classes of models up to information equivariance. For example, the number of symmetries that render two models equivalent increases vastly in the number of latent variables considered. These information invariances are not generally factored into current methods, which means that the model space they face is overwhelmingly larger than it need be. Quotienting by these (and other) invariances should greatly reduce the complexity of the problem—and highly improve model evidence. Furthermore, specifying the prior over equivalence classes has an advantage. Consider that the prior probability of a model equivalence class equals the sum of the prior probabilities for all elements in the class. When specifying prior probabilities over individual models as opposed to equivalence classes, we may have unintended over-counting effects; leading to a higher prior for model equivalence classes with a large number of elements, such as those comprised of models with a high number of latent variables. Practically addressing this issue seems important to scale Bayesian structure learning.
@@ 4.3.1 Core knowledge priors
Core knowledge therefore constitutes knowledge that is valid across any naturalistic world, which translates into significant constraints on the prior over models. The prior over models as explanations for the world is then constrained by the consistency with the underlying information geometry (i.e. local constraints) and core knowledge constraints (i.e. non-local constraints). [^4] The maximal entropy probability distribution that is consistent with an underlying geometry and some constraints is a Gibbs measure on the underlying metric space [94, Chapter 9].
@@ 5 AI alignment
In what follows, we approach AI alignment through the lens of active inference and structure learning, using Asimov’s Three Laws of robotics [118] as a simple illustrative example. Our focus is on providing new ways of thinking about the AI alignment problem, rather than recommending a specific solution to be implemented. The main conceptual point is that we can frame alignment as taking actions that comply with the other’s preferences, and we can infer those preferences through structure learning, which, in the language of psychology, corresponds to instantiating a sophisticated form of theory of mind. What follows is necessarily conceptual, as a practical implementation would first require structure learning active inference agents as presented in Section 4 that have the capability to represent other agents with different generative models—a capability that remains to be developed.
@@ 5.1 Well-being, alignment and cautious AI
where $d\triangleq d(t)$ is the data entertained by the agent at time $t$; necessarily a subset of past and present states of the agent (boundary and internal states). Note that well-being, when expressed in this way, is quantified in terms of natural units of information (nats). This definition of well-being is fairly established in the active inference literature [119, 120, 121, 122]. Additionally, when harm is quantified in this way, the resulting equations of motion from active inference can reproduce well-known empirical phenomena observed in biological collectives under harm [123], lending additional (post-hoc) validity to this definition.
@@
Beyond these conceptual points, active inference also offers algorithmic advances for building safer and more robust AI systems [19, 18].
@@ 5.2 Toward empathetic AI
We now define an empathetic agent as an agent that not only models the objects in its external world, but also the other agents within it as well as their mental (i.e. internal) states. This effectively recognizes that the external world comprises other agents’ actions, observations and internal states, which may be inferred from data, along with all other external states in the shared environment. Technically, an agent that has beliefs about another’s mental state is capable of empathy. In the language of cognitive psychology, empathetic AIs have theory of mind.
Instantiating this in practice is a difficult problem, and may require entertaining generative world models as in Figure 10 (left panel). An empathetic agent may, in principle, infer another agent’s world model from the other’s inferred observations and actions (assuming, for instance, that it takes actions that minimize expected free energy—or any other objective—and solving the inverse problem). Note that assuming that the other possesses the same generative model as oneself ("you are like me") greatly simplifies the process by which one forms beliefs about another, because my model of me becomes my model of you [125].
@@
However, it is crucial to recognize that empathy alone does not guarantee benevolence or safety. Competitive or even adversarial agents may benefit from sophisticated models of others for instrumental reasons. In particular they may use these to manipulate, or deceive others to achieve their goals [128, 129]. The key might be to prescribe agent’s actions to be benevolent in the sense of complying with the other’s preferences. However, note that benevolent agents may still choose to deceive, as it may happen to be beneficial to manipulate humans in the process of optimizing their well-being (cf. white lies).
@@ 5.3 The Three Laws of Robotics
To mathematically instantiate the First Law, first consider an empathetic agent whose sole preferences are to prevent another agent (the target agent) from coming to harm. The empathetic agent’s actions could be governed by the expected free energy (4) where we discard all external variables besides the target agent’s harm
@@
Here FE is the harm (or negative well-being) of the target agent—the only external variable that is retained in the expected free energy (4) which governs the empathetic agent’s actions.
A version of the First Law could be achieved by assuming a preferred distribution over the target agent’s harm in (11), e.g. an exponentially decaying probability distribution over harm, so that zero or low harm would exponentially preferred to high harm, or by (soft-)thresholding the target’s well-being to desirable values. [^8] In any case, the fact that the preferred distribution in (11) depends on the history of data means that preferences can be inferred, or learned, over time in a data-driven way [42, 40]. To appropriately model the target’s level of harm, the empathetic agent would have to model many other external variables; however the only imperatives driving its actions would be to resolve ambiguity—and realize the preferred distribution—about the target’s well-being.
@@ 5.4 Alignment beyond perspective taking
Numerical studies suggest that populations may naturally converge to free energy equilibria, where all agents share the same generative model (and preferences), albeit possibly at an evolutionary timescale [133, 134]. In this case, each agent’s goal is the group’s (shared) goal and empathy is an emergent property [125]. Clearly, these numerical studies need to be extended to empathetic agents and agents that learn their own preferences [42, 40].
@@
5.5 Related work
Related work uses causal models of agent-environment interactions, but with the aim of analysing agent’s incentives [131, 130]. This provides a complementary perspective on analyzing and designing AI systems with desirable incentive structures, and aligns well with our discussion of theory of mind, including its potential for benevolent as well as adversarial uses. The structure learning approach we develop may provide a way to dynamically construct these causal networks and go beyond identifying the mere presence or absence of various properties such as incentives [131], intentions [135], and deception [136], by quantifying these phenomena.
Our work builds upon Bayesian models of theory of mind. Classical work demonstrates Bayesian inference of others’ mental states and reward functions [138], and that empathy can be straightforward when another has the same generative model as oneself [125]. Recent work explores multi-agent cooperation through utility function inference [139], but maintain only a fixed repertoire of agent profiles for classification. In all cases, approaches assume known generative model structures, whereas we emphasize structure learning about agents whose generative model is unknown. Furthermore, in contrast to previous work, we emphasise the importance of epistemic actions (e.g. asking questions) to gain information about other agents’ generative model and preferences.
@@ 6 Discussion
Active inference can be arrived at from two opposite directions: the bottom-up, inductive and historical approach, which is where theories of brain function were continuously refined and generalized to account for various empirical phenomena [5, 6, 8, 9]. A top-down, deductive approach, through the nascent field of physics called Bayesian mechanics [140, 33, 10], which bridges elementary descriptions of particles, things and agents in the physical world, with descriptions of inference. From a theoretical perspective, there are many opportunities to further develop this top-down approach, e.g. 1) by characterizing the subclasses of physical systems that exhibit higher-order cognitive phenomena [32, 141, 142], and 2) by making these formulations mathematically rigorous, using purposefully developed tools in the theory of stochastic processes [34].
@@ Acknowledgements
A.R. is funded by the Australian Research Council (Ref: DP200100757). A.R. is also funded by Australian National Health and Medical Research Council Investigator Grant (Ref: 1194910). A.R. is affiliated with The Wellcome Centre for Human Neuroimaging supported by core funding from Wellcome [203147/Z/16/Z]. A.R. is a CIFAR Azrieli Global Scholar in the Brain, Mind & Consciousness Programme.
Competing interests
The authors declare no competing interests.